OpenStudy (anonymous):

math section down someone help f(x)= x^3-2x^2 relative min and max using second derivative test?

hero (hero):

You only need to post your question once. The math section is not down.

OpenStudy (anonymous):

im sorry can you link me to it? it wasnt work for me.

hero (hero):

Aren't you already in the math section bro?

OpenStudy (anonymous):

if you know how to do this, can you help me? plsss

OpenStudy (anonymous):

they split them into subsections? before algebra,calc,etc were all in one sections

OpenStudy (anonymous):

pls can you help me wi

OpenStudy (anonymous):

*with this

OpenStudy (anonymous):

im preparing for a test plssss

hero (hero):

1. Take the derivative 2. Set it to zero, solve for x 3. Take second derivative 4. Input value of x 5. If negative, then x is max, If positive, then x is min

OpenStudy (anonymous):

ok so far ive got f`(x)=3x^2-4x 3x^2-4x=0 x(3x-4)=0 whats next?

hero (hero):

Take the derivative of f'(x) = 3x^2 - 4x to find f''(x) To find the zeroes of f'(x) use the zero product property

OpenStudy (anonymous):

im confused...can you demonstrate it pls?

hero (hero):

Pretend that f'(x) is your new h(x) Now take the derivative of h(x) = 3x^2 - 4x

OpenStudy (anonymous):

6x-4 f"(x)

OpenStudy (anonymous):

??

hero (hero):

Okay good

hero (hero):

We need to input the zeroes of f'(x) into f''(x)

hero (hero):

I'm pretty sure you know how to solve x(3x - 4) = 0

OpenStudy (anonymous):

x=0, x=4/3 ??

hero (hero):

exactly, now put those values into f''(x) one at a time.

OpenStudy (anonymous):

6(0)-4= -4 6(4/3)-4= 4

OpenStudy (anonymous):

??

hero (hero):

Okay, so which is min and which is max?

OpenStudy (anonymous):

min=-4 max=4

hero (hero):

No, you have it backwards. Re-read the steps I presented to you.

OpenStudy (anonymous):

oh max=-4 min=4

OpenStudy (anonymous):

thanks i have another one f(x)=x^3 + 3x^2 but let me attempt it first

hero (hero):

That's the only thing about second derivative. You have to remember that if f''(x) produces negative value, then f(x) is max, but if f''(x) produces positive value, then f(x) is min

hero (hero):

Yes, you should try it first.

hero (hero):

No reason why you can't finish it. The only thing is you need to find f(0) and f(4/3) for the previous problem.

OpenStudy (anonymous):

ive got...

hero (hero):

If I were you, I would make a chart to keep track of everything.

OpenStudy (anonymous):

f'(x)=3x^2+6x=0 3x(x+2)=0 x=0 x=2 f"(x)=6x + 6) 6(0)+6=6 6(2)+6=18

hero (hero):

You mean x = -2

OpenStudy (anonymous):

how so

hero (hero):

Take a moment to think about it

OpenStudy (anonymous):

hey the program went off are you there

OpenStudy (anonymous):

i figured out

hero (hero):

Okay good

OpenStudy (anonymous):

i have a quite similar question...now i have to find the absolute max and min of f(x)= x^3 + 3x^2 on the interval [1,1]...whats the difference between relative and absolute anyhow

hero (hero):

For the first couple problems, there was only one min and one max. Now for this, you will have to choose the highest min and max between several values.

hero (hero):

To do this, you do the same exact thing you did before, except this time, you will also have to test x = 1 and x = -1

OpenStudy (anonymous):

so i take the first and second derivatives then test -1 and 1? into which derivative?

hero (hero):

Basically, do everything you did before (all five steps while ignoring x = -1,1 for the moment). Then after you have completed the five steps, then also evaluate x = -1,1

OpenStudy (anonymous):

might need a demonstration...its kind of confusing..let me attempt first

hero (hero):

Basically just pretend you don't have [1,1] for now.

hero (hero):

Pretend that you have to find the relative min and max of f(x)= x^3 + 3x^2

hero (hero):

So perform the five steps as usual.

hero (hero):

Let me know what you get afterwards

OpenStudy (anonymous):

f'(x)=3x^2+6x x(3x-6)=0 x=0 x=6/3 f"(x)=6x+6 6(0)+6=6 6(6/3)+6=

OpenStudy (anonymous):

??

hero (hero):

6/3 = 2 btw

OpenStudy (anonymous):

x^3 + 3x^2 was orginal function

OpenStudy (anonymous):

yea i forgot to type 18

hero (hero):

What? Where does 18 come from? I said 6/3 = 2

OpenStudy (anonymous):

2 * 6 + 6=18?

hero (hero):

Okay, but

hero (hero):

We only need to consider x values between -1 and 1 so x = 2 is out

OpenStudy (anonymous):

ohhh ok

hero (hero):

So basically you need to test: f''(-1) f''(0) f''(1)

OpenStudy (anonymous):

ohh gotcha

hero (hero):

If I were you, I would just simply test f(-1) f(0) f(1) That will tell you the min and max

hero (hero):

You're only looking for the extreme min and extreme max

OpenStudy (anonymous):

i see thanks for the help. i have several more questions though

hero (hero):

I bet you do

OpenStudy (anonymous):

how do i solve 3^(x^2 - x) = (1/9)^(3x)

OpenStudy (anonymous):

im preparing for a test sorry

hero (hero):

For this, I'll need to write out the steps so that you can see them

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

can you pls?

hero (hero):

Hang on a sec bro

OpenStudy (anonymous):

ok

hero (hero):

Okay, I have it

hero (hero):

I'll send you a link to vyew for this

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

is that the final answer

hero (hero):

No, bro. It is freezing up

OpenStudy (anonymous):

ohhh ok

hero (hero):

Good luck figuring out the steps

hero (hero):

You probably got lost somewhere. I just know it

OpenStudy (anonymous):

yeah bro ill try...you think you can help with this last one? 3+ 2e^(-t)= 7

hero (hero):

That's an easy one bro

OpenStudy (anonymous):

have to solve the equation....ill try to go over it and ask my peers for help

OpenStudy (anonymous):

damn bro im so tire from all this studyy...i really cant figure out that one

hero (hero):

Go back to vyew bro

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for all the help bro